OpenStudy (donnie1999):

HELP ME PLEASE. Solve the equation x^2 − 36 = 0.

OpenStudy (anonymous):

First we need to isolate the x^2 so we add 36 to both sides of the equation. What will you end up with?

Mehek (mehek14):

\(x^2 = 0 + 36\)

Mehek (mehek14):

what's 0+36

Mehek (mehek14):

@donnie1999 ?

OpenStudy (anonymous):

So you have x^2 = 36 correct?

OpenStudy (donnie1999):

soorry

OpenStudy (anonymous):

What happens if you take the square root of both sides?

Mehek (mehek14):

\(x=\sqrt{36}\)

OpenStudy (anonymous):

and the square root of 36 is...

OpenStudy (donnie1999):

6